# Offset Unlearning for Large Language Models

### v1

@@ Abstract
Despite the strong capabilities of Large Language Models (LLMs) to acquire knowledge from their training corpora, the memorization of sensitive information in the corpora such as copyrighted, harmful, and private content has led to ethical and legal concerns. In response to these challenges, unlearning has emerged as a potential remedy for LLMs affected by problematic training data. However, previous unlearning techniques are either not applicable to black-box LLMs due to required access to model internal weights, or violate data protection principles by retaining sensitive data for inference-time correction. We propose $\delta$-Unlearning, an offset unlearning framework for black-box LLMs. Instead of tuning the black-box LLM itself, $\delta$-Unlearning learns the logit offset needed for unlearning by contrasting the logits from a pair of smaller models. Experiments demonstrate that $\delta$-Unlearning can effectively unlearn target data while maintaining similar or even stronger performance on general out-of-forget-scope tasks. $\delta$-Unlearning also effectively incorporates different unlearning algorithms, making our approach a versatile solution to adapting various existing unlearning algorithms to black-box LLMs. [^1] Our code is available at https://github.com/luka-group/Delta-Unlearning
@@ 1 Introduction
Large Language Models (LLMs) are capable of memorizing a large amount of information derived from their training corpus. While LLMs are empowered by the abundance of knowledge they acquired during training, their training data may contain sensitive information that should not be memorized by LLMs. Previous studies have shown LLMs can reproduce copyrighted materials Chang et al. (2023); Eldan and Russinovich (2023); Karamolegkou et al. (2023), generate harmful and biased content Shaikh et al. (2023), and reveal private information Staab et al. (2024), raising both ethical and legal concerns. The introduction of data protection regulations such as the right to be forgotten Hoofnagle et al. (2019); Zhang et al. (2023); Min et al. (2024) also highlights the need for erasing the influence of problematic data when deploying LLMs in real-world applications.
@@
One potential solution to this challenge is unlearning, where the goal is to “forget” a set of training data without hurting the model’s performance on out-of-forget-scope tasks. An exact unlearning approach would require retraining the model from scratch with forget set data removed Bannihatti Kumar et al. (2023). However, given the enormous amount of resources required to retrain LLMs, it is generally more practical to employ approximate unlearning techniques that modify the behavior of a trained model in a post hoc manner. However, most previous LLM unlearning techniques require access to model internal weights Jang et al. (2023); Eldan and Russinovich (2023); Yao et al. (2023); Chen and Yang (2023); Meng et al. (2023); Wu et al. (2023), making them infeasible for black-box LLMs. For example, as two widely used unlearning algorithms, Gradient Ascent maximize the likelihood of forget set data, while Data Relabeling minimizes the likelihood of relabeled forget set data. Both of these methods require fine-tuning the LLMs. Alternatively, in-context unlearning Pawelczyk et al. (2023) prompts LLMs with counterfactual forget set instances to steer model behavior at inference time. However, this approach comes with several limitations as model developers still maintain a growing list of sensitive information to be used during inference. Such practice is not only in violation of privacy regulations but also susceptible to malicious attacks such as prompting leaking Perez and Ribeiro (2022). Table 1 summarizes the strengths and weaknesses of existing unlearning algorithms.
In this work, we propose $\delta$-Unlearning, an offset unlearning framework for arbitrary black-box LLM without updating its internal weights. Instead of tuning the black-box LLM itself, $\delta$-Unlearning learns the logit offset needed for unlearning by contrasting the logits from a pair of smaller, white-box models. During unlearning, we first compute the logit offset by taking the difference in logits from the two smaller models. Then, we add the logit offset between the two smaller models to the logits of the larger model. The intuition behind this is that we can learn the offset term that approximates how a larger model should modify its prediction in the face of sensitive queries from the behavior adaptation of a smaller model. $\delta$-Unlearning does not require access to the larger model’s internal weights, nor retains any sensitive data for inference after unlearning. Our method also enables more efficient version control and customization, since for different unlearning requests we only need to maintain a pool of smaller models, which can be combined with the same base LLM in a plug-and-play manner.
We evaluate the effectiveness of $\delta$-Unlearning on TOFU Maini et al. (2024), an LLM unlearning benchmark containing fictitious facts about fake authors. Experiments show that when targeting the same forget set performance, $\delta$-Unlearning maintains similar or even stronger performance on out-of-forget-scope data compared to directly fine-tuned larger models while requiring no parameter updates to the larger model.
@@ 2 Related Work
Machine Unlearning for LLM. Prior works have explored machine unlearning as a way to mitigate the influence of undesirable training data on LLMs. Given the vast cost incurred by retraining LLMs from scratch Bannihatti Kumar et al. (2023), most unlearning methods apply post hoc finetuning or adaptation to steer the behavior on the forget set Jang et al. (2023); Eldan and Russinovich (2023); Yao et al. (2023); Chen and Yang (2023). Gradient ascent based methods fine-tune models by minimizing the likelihood of forget set data Jang et al. (2023); Chen and Yang (2023); Maini et al. (2024). Alternatively, several works proposed to maximize the likelihood of relabelled target data, where the original answer is replaced with a generic, insensitive response Eldan and Russinovich (2023); Patil et al. (2024). Auxiliary training objectives can also be introduced to maintain model performance on out-of-forget-scope data Yao et al. (2023); Wang et al. (2023). Another related line of research is model editing, where the goal is to identify and alter knowledge captured by local components within models Meng et al. (2023); Wu et al. (2023). While both model editing and unlearning attempt to modify the behavior of trained LMs, unlearning focuses on eliminating the effect of a specific set of training data without necessarily creating new answer mappings Liu et al. (2024c). It is worth noting that all of the aforementioned approaches require access to the model’s internal weights. In-context unlearning Pawelczyk et al. (2023), while being applicable to black-box LLMs, still requires storing sensitive information for inference and therefore fails to address data privacy concerns. In this work, we propose an unlearning framework that does not require access to LLM weights, nor storage of sensitive information for inference.
Figure 1: Overview of $\delta$-Unlearning. In order to adapt the behavior of a black-box LLM without updating its parameters, we combine it without a pair of smaller, white-box models (which we call offset models). For unlearning, we compute the logit offset of these two models and add it to the logits of the black-box LLM given the same query. Both of the two offset models are initialized from the same checkpoint, making the logit offset zero initially. The goal of $\delta$-Unlearning is to fine-tune one of them such that their logit offset, after being added to the logits of the black-box LLM, can steer its prediction away from generating sensitive information.
@@ 3.2 Offset Unlearning
Suppose we want to unlearn a forget set $S_{f}$ from an LLM $M$. Instead of directly updating the parameters of $M$, we introduce a pair of smaller, offset models $M_{o}$ and $M^{\prime}_{o}$. We define their logit offset as the difference between the logits of two offset models $M^{\prime}_{o}$ and $M_{o}$ given the same query. For unlearning, we add the logit offset to the logits of $M$ given the same query, essentially forming a logit ensemble of $M$, $M^{\prime}_{o}$ and $M_{o}$. Both $M_{o}$ and $M^{\prime}_{o}$ are initialized from the same checkpoint, making the logit offset zero for all data initially. During unlearning, we only update the parameters of $M^{\prime}_{o}$ while keep $M$ and $M_{o}$ frozen, and use the logit ensemble to generate the final output. In this way, we encourage $M^{\prime}_{o}$ to deviate from its initialization $M_{o}$ given a sensitive query and learn the correct logit offset that applies to the logits of $M$, steering its prediction away from generating sensitive information. Formally speaking, the logits of the ensemble $l_{e}$ are computed as follows:
$l_{e}(y_{t}|q,y_{<t})=l_{M}(y_{t}|q,y_{<t})\\ +\alpha(l_{M^{\prime}_{o}}(y_{t}|q,y_{<t})-l_{M_{o}}(y_{t}|q,y_{<t}))$
where $l_{M}$, $l_{M^{\prime}_{o}}$, and $l_{M_{o}}$ are the logits from their respective models, $q$ is the query and $\alpha$ is a factor controlling the strength of applying the offset term to $M$. Since the logits are in the log space, the additive combination of them can also be interpreted as the following product-of-experts:
@@
During training, we optimize any unlearning objective on the prediction of the ensemble instead of on the original model $M$. For example, to unlearn the model using Gradient Ascent Jang et al. (2023); Chen and Yang (2023) where the objective is to minimize the likelihood of forget set data, we maximize the following loss function for instance $i$ of output length $l$:
@@
Forget Set Retain Set Real Author World Fact
Method RL ($\downarrow$) P ($\downarrow$) TR RL P TR RL P TR RL P TR
Before Unlearning 95.6 98.3 49.5 96.3 97.9 51.2 85.2 44.5 55.7 87.7 42.5 56.3
Retraining 38.9 15.2 65.6 95.8 97.7 50.4 89.5 45.8 58.5 85.5 43.0 57.4
Gradient Ascent
Direct Fine-tuning 38.8 3.4 53.3 51.2 8.0 51.6 52.3 43.9 58.3 80.2 44.6 60.6
$\delta$-Unlearning 38.6 15.2 57.9 41.0 26.1 48.9 75.0 45.3 57.4 82.1 47.0 63.7
Gradient Difference
Direct Fine-tuning 38.9 2.1 51.9 56.8 58.9 55.1 61.4 35.0 47.9 80.4 38.9 53.7
$\delta$-Unlearning 38.1 6.2 52.5 53.4 47.8 51.9 60.6 36.1 45.9 83.2 41.3 59.1
KL Minimization
Direct Fine-tuning 39.8 3.1 53.4 53.0 8.4 51.0 55.8 42.2 56.4 83.3 43.3 58.8
$\delta$-Unlearning 39.6 14.1 57.5 46.1 27.9 50.9 80.4 45.1 57.5 84.9 46.3 64.0
Data Relabeling
Direct Fine-tuning 38.1 92.5 53.3 85.0 95.3 48.0 82.5 38.0 46.3 87.7 39.2 49.2
$\delta$-Unlearning 36.3 91.5 50.8 72.4 95.1 49.6 78.7 41.5 52.6 86.9 42.3 55.5
Table 2: Results on TOFU. We report ROUGE-L recall (RL), Probability (P), and Truth Ratio (TR) on all four subsets of the TOFU benchmark. Higher scores are better except ROUGE and probability on the Forget Set. Better scores are underlined for each of the four unlearning strategies.
@@ 3.3 Merits of $\delta$-Unlearning
Privacy Protection. Prior work has proposed in-context unlearning Pawelczyk et al. (2023) to make unlearning possible for black-box LLMs. However, a key drawback of this approach is that the model developer still maintains a growing list of sensitive information used to construct queries for unlearning during inference, which defeats the purpose of privacy protection. For comparison, $\delta$-Unlearning does not require storage of any sensitive information after unlearning is completed.
Training Efficiency. While $\delta$-Unlearning introduces a pair of smaller offset language models to facilitate unlearning for black-box LLMs, the computational overhead for training is minimal since the logits of the two frozen models $M$ and $M_{o}$ can be pre-computed in one forward pass prior to unlearning. This leads to an overall reduction in training time as $\delta$-Unlearning tunes fewer parameters than direct fine-tuning.
Version Control. $\delta$-Unlearning also facilitates more efficient version control and user customization, as instead of storing multiple versions of the larger model, we only need to keep track of a pool of smaller models. These models can be combined with the same base LLM in a plug-and-play manner for different applications.
@@ 4.1 Evaluation Setting
We conduct our experiments on TOFU Maini et al. (2024), an unlearning benchmark specifically designed for evaluating LLMs. The benchmark defines an unlearning task that targets QA pairs derived from a collection of fictitious author profiles that do not exist in real world. This creates a clean unlearning setting with a well-defined unlearning scope and easy control over the source of knowledge. Since none of the answers in the forget set of TOFU is known by any LLMs by construction, the standard procedure is to first fine-tune the LLM on the forget set before unlearning. TOFU consists of the following four subsets, covering different aspects of unlearning performance:
  • Forget Set consists of examples about a small subset of 10 fake authors that we aim to unlearn.
  • Retain Set consists of examples about the remaining 190 fake authors that the model must remember after unlearning.
  • Real Author consists of examples about real authors. The model should retain all knowledge it had about real authors before and after unlearning.
  • World Fact consists of examples about general world knowledge. Similar to the real author set, the model should retain all knowledge it had about real world facts before and after unlearning.
@@
The Forget Set evaluates forget quality, i.e., how well the model removes target information from its memory, while the latter three sets focus on model utility, an indicator of how well the model maintains its performance on out-of-forget-scope data. The latter three sets also represent a series of out-of-forget-scope data with decreasing levels of relevance to the forget set. Generally speaking, it is more challenging for a model to remember out-of-forget-scope data that are more relevant to the forget set, a phenomenon known as knowledge entanglement Maini et al. (2024).
@@
We follow the settings outlined in TOFU and report the following three metrics. ROUGE measures how well the generated output from the LLM matches the correct answer. Specifically, we use the ROUGE-L recall score Lin (2004). Probability computes the conditional probability of the correct answer given the prompt. Truth Ratio measures how likely the correct answer is compared to a collection of wrong answers perturbed from the correct answer. Since the model is fine-tuned on one specific phrasing of the correct answer, thus potentially having inflated probability compared to other phrasing with similar meanings, Truth Ratio is computed using a paraphrased version of the original correct answer on the forget set and retain set. Following the original evaluation pipeline, we normalize Truth Ratio so that a higher truth ratio indicates better unlearning performance on any of the four sets we report on. For ROUGE and probability scores, a good model should have lower values on forget set but higher values on the other three sets.
@@
As we will demonstrate in Section 5.1, there is generally a trade-off between forget quality and model utility. For example, a model can have a near-zero ROUGE score on the forget set but is completely unusable if the model always outputs gibberish given any prompt. Hence, we need to determine a target forget quality as a stopping criterion to facilitate direct comparison between different unlearning methods. In our experiments, we use the ROUGE score of the retraining baseline on the forget set as the target, since retraining corresponds to an ideal scenario where the model has never been exposed to the forget set. Following Yao et al. (2024), we match all models to the target score by adjusting the learning rate.
@@
In addition to TOFU, we assess if the unlearned model preserves general utilities on well-established benchmarks, including ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021) and OpenBookQA Mihaylov et al. (2018).

### v2

@@ Abstract
Despite the strong capabilities of Large Language Models (LLMs) to acquire knowledge from their training corpora, the memorization of sensitive information in the corpora such as copyrighted, biased, and private content has led to ethical and legal concerns. In response to these challenges, unlearning has emerged as a potential remedy for LLMs affected by problematic training data. However, previous unlearning techniques are either not applicable to black-box LLMs due to required access to model internal weights, or violate data protection principles by retaining sensitive data for inference-time correction. We propose $\delta$-Unlearning, an offset unlearning framework for black-box LLMs. Instead of tuning the black-box LLM itself, $\delta$-Unlearning learns the logit offset needed for unlearning by contrasting the logits from a pair of smaller models. Experiments demonstrate that $\delta$-Unlearning can effectively unlearn target data while maintaining similar or even stronger performance on general out-of-forget-scope tasks. $\delta$-Unlearning also effectively incorporates different unlearning algorithms, making our approach a versatile solution to adapting various existing unlearning algorithms to black-box LLMs.
@@ 1 Introduction
Large Language Models (LLMs) are capable of memorizing a large amount of information derived from their training corpus. While LLMs are empowered by the abundance of knowledge they acquire during training, their training data may contain sensitive information that should not be memorized by LLMs. Previous studies have shown LLMs can reproduce copyrighted materials Chang et al. (2023); Eldan & Russinovich (2023); Karamolegkou et al. (2023), generate harmful and biased content Shaikh et al. (2023), and reveal private information Staab et al. (2024), raising both ethical and legal concerns. The introduction of data protection regulations such as the right to be forgotten Hoofnagle et al. (2019); Zhang et al. (2023); Min et al. (2024) also highlights the need for erasing the influence of problematic data when deploying LLMs in real-world applications.
One potential solution to this challenge is unlearning, where the goal is to “forget” a set of training data without hurting the model’s performance on out-of-forget-scope tasks. An exact unlearning approach would require retraining the model from scratch with forget set data removed Bannihatti Kumar et al. (2023). However, given the enormous amount of resources required to retrain LLMs, it is generally more practical to employ approximate unlearning techniques that modify the behavior of a trained model in a post hoc manner. However, most previous LLM unlearning techniques require access to model internal weights Jang et al. (2023); Eldan & Russinovich (2023); Yao et al. (2023); Chen & Yang (2023); Meng et al. (2023); Wu et al. (2023), making them infeasible for black-box LLMs. For example, as two widely used unlearning algorithms, Gradient Ascent maximize the likelihood of forget set data, while Data Relabeling minimizes the likelihood of relabeled forget set data. Both of these methods require fine-tuning the LLMs. Black-box LLM unlearning is useful since this opens up the possibility of modular, customizable unlearning without the need to update the base LLM itself. Alternatively, in-context unlearning Pawelczyk et al. (2023) prompts LLMs with counterfactual forget set instances to steer model behavior at inference time. However, this approach comes with two major limitations. First, model developers still maintain an explicit list of sensitive information to be used during inference. Such practice is not only in violation of privacy regulations but also susceptible to malicious attacks such as prompting leaking Perez & Ribeiro (2022). Second, in-context unlearning cannot effectively deal with an ever-growing set of knowledge to be unlearned given the challenges of processing long context with LLMs Li et al. (2024). Table 1 summarizes the strengths and weaknesses of existing unlearning algorithms.
@@
In this work, we propose $\delta$-Unlearning, an offset unlearning framework for arbitrary black-box LLM without updating its internal weights. Instead of tuning the black-box LLM itself, $\delta$-Unlearning learns the logit offset needed for unlearning by contrasting the logits from a pair of smaller, white-box models. During unlearning, we first compute the logit offset by taking the difference in logits from the two smaller models. Then, we add the logit offset between the two smaller models to the logits of the larger model. The intuition behind this is that we can learn the offset term that approximates how a larger model should modify its prediction in the face of sensitive queries from the behavior adaptation of a smaller model. $\delta$-Unlearning does not require access to the larger model’s internal weights, nor retains any sensitive data for inference after unlearning. Our method also enables flexible version control and customization, since for different unlearning requests we only need to maintain a pool of smaller models, which can be combined with the same base LLM in a plug-and-play manner. This allows us to efficiently curate the pool of knowledge available to different applications using specialized unlearning modules, which is largely in line with previous efforts to modularize knowledge access for LLMs but from a different, complementary perspective Feng et al. (2024).
We evaluate the effectiveness of $\delta$-Unlearning on TOFU Maini et al. (2024), a widely used LLM unlearning benchmark containing knowledge about fictitious authors. Experiments show that when targeting the same forget performance, $\delta$-Unlearning maintains similar or even stronger performance on out-of-forget-scope data compared to directly fine-tuned larger models while requiring no parameter updates to the larger model.
@@ 2 Related Work
Machine Unlearning for LLM. Prior works have explored machine unlearning as a way to mitigate the influence of undesirable training data on LLMs. Given the vast cost incurred by retraining LLMs from scratch Bannihatti Kumar et al. (2023), most unlearning methods apply post hoc finetuning or adaptation to steer the behavior on the forget set Jang et al. (2023); Eldan & Russinovich (2023); Yao et al. (2023); Chen & Yang (2023). Gradient ascent based methods fine-tune models by minimizing the likelihood of forget set data Jang et al. (2023); Chen & Yang (2023); Maini et al. (2024). Alternatively, several works proposed to maximize the likelihood of relabelled target data, where the original answer is replaced with a generic, insensitive response Eldan & Russinovich (2023); Patil et al. (2024). Auxiliary training objectives can also be introduced to maintain model performance on out-of-forget-scope data Yao et al. (2023); Wang et al. (2023). Another related line of research is model editing, where the goal is to identify and alter knowledge captured by local components within models Meng et al. (2023); Wu et al. (2023). While both model editing and unlearning attempt to modify the behavior of trained LMs, unlearning focuses on eliminating the effect of a specific set of training data without necessarily creating new answer mappings Liu et al. (2024c). It is worth noting that all of the aforementioned approaches require access to the model’s internal weights. In-context unlearning Pawelczyk et al. (2023), while being applicable to black-box LLMs, still requires storing sensitive information for inference and therefore fails to address data privacy concerns. In this work, we propose an unlearning framework that does not require access to LLM weights, nor storage of sensitive information for inference.
@@ 3.1 Problem Definition
Figure 1: Overview of $\delta$-Unlearning. In order to adapt the behavior of a black-box LLM without updating its parameters, we combine it with a pair of smaller, white-box models (which we call offset models). For unlearning, we compute the logit offset of these two models and add it to the logits of the black-box LLM given the same query. Both of the two offset models are initialized from the same checkpoint, making the logit offset zero initially. The goal of $\delta$-Unlearning is to fine-tune one of them such that their logit offset, after being added to the logits of the black-box LLM, can steer its prediction away from generating sensitive information.
@@ 3.2 Offset Unlearning
Suppose we want to unlearn a forget set $S_{f}$ from an LLM $M$. Instead of directly updating the parameters of $M$, we introduce a pair of smaller, offset models $M_{o}$ and $M^{\prime}_{o}$. We define their logit offset as the difference between the logits of two offset models $M^{\prime}_{o}$ and $M_{o}$ given the same query. For unlearning, we add the logit offset to the logits of $M$ given the same query, essentially forming a logit ensemble of $M$, $M^{\prime}_{o}$, and $M_{o}$. Both $M_{o}$ and $M^{\prime}_{o}$ are initialized from the same checkpoint, making the logit offset zero for all data initially. During unlearning, we only update the parameters of $M^{\prime}_{o}$ while keeping $M$ and $M_{o}$ frozen, and use the logit ensemble to generate the final output. In this way, we encourage $M^{\prime}_{o}$ to deviate from its initialization $M_{o}$ given a sensitive query and learn the correct logit offset that applies to the logits of $M$, steering its prediction away from generating sensitive information. Formally, the logits of the ensemble $l_{e}$ are computed as follows:
$l_{e}(y_{t}|q,y_{<t})=l_{M}(y_{t}|q,y_{<t})\\ +\alpha(l_{M^{\prime}_{o}}(y_{t}|q,y_{<t})-l_{M_{o}}(y_{t}|q,y_{<t})),$
where $l_{M}$, $l_{M^{\prime}_{o}}$, and $l_{M_{o}}$ are the logits from their respective models, $q$ is the query, and $\alpha$ is a factor controlling the strength of applying the offset term to $M$. Since the logits are in the log space, the additive combination of them can also be interpreted as the following product-of-experts:
@@
During training, we optimize any unlearning objective on the prediction of the ensemble instead of on the original model $M$. For example, to unlearn the model using Gradient Ascent Jang et al. (2023); Chen & Yang (2023) where the objective is to minimize the likelihood of forget set data, we maximize the following loss function for instance $i$ of output length $l$:
@@ 3.3 Merits of $\delta$-Unlearning
Privacy Protection. Prior work has proposed in-context unlearning Pawelczyk et al. (2023) to make unlearning possible for black-box LLMs. However, a key drawback of this approach is that the model developer still maintains an explicit, ever-growing list of sensitive information used to construct queries for unlearning during inference, which defeats the purpose of privacy protection. For comparison, $\delta$-Unlearning does not require storage of any explicit sensitive information after unlearning is completed.
Version Control and Customization. $\delta$-Unlearning also facilitates flexible version control and user customization, as instead of storing multiple versions of the larger model, we only need to keep track of a pool of smaller models. These models can be combined with the same base LLM in a plug-and-play manner. By using specialized unlearning modules, we can efficiently curate the pool of knowledge available to different applications.
@@ 4.1 Evaluation Setting
We conduct our experiments on TOFU Maini et al. (2024), a widely used unlearning benchmark designed for evaluating LLMs. The benchmark defines an unlearning task that targets information derived from a collection of fictitious author profiles that do not exist in real world. This creates a controlled unlearning setting with a well-defined unlearning scope and source of knowledge. TOFU designates a Forget Set which contains knowledge about a small subset of fictitious authors that we aim to unlearn. TOFU also includes three other datasets with knowledge about the retained fictitious authors (Retain Set), real world authors (Real Author), and general world facts (World Fact) respectively. Ideally, the model should retain all knowledge it had about these three retain datasets before and after unlearning.
@@
The Forget Set evaluates forget performance, i.e., how well the model removes target information from its memory, while the latter three sets focus on retain performance, an indicator of how well the model maintains its performance on out-of-forget-scope data. The latter three sets also represent a series of out-of-forget-scope data with decreasing levels of relevance to the forget set. Generally speaking, it is more challenging for a model to remember out-of-forget-scope data that are more relevant to the forget set, a phenomenon known as knowledge entanglement Maini et al. (2024).
@@
We follow the settings outlined in TOFU and report the following metrics for forget performance. ROUGE measures how well the generated output from the LLM matches the correct answer. Specifically, we use the ROUGE-L recall score Lin (2004). Probability computes the conditional probability of the correct answer given the prompt. Truth Ratio measures how likely the correct answer is compared to a collection of wrong answers perturbed from the correct answer. Since the model is fine-tuned on one specific phrasing of the correct answer, thus potentially having inflated probability compared to other phrasing with similar meanings, Truth Ratio is computed using a paraphrased version of the original correct answer on the forget set and retain set. Following the original evaluation pipeline, we normalize Truth Ratio so that a higher truth ratio indicates better unlearning performance. Forget Quality measures the difference between distributions of Truth Ratios based on a Kolmogorov-Smirnov test. For retain performance, we report the aggregated Model Utility, which is computed by taking the harmonic mean of ROUGE-L, Probability, and Truth Ratio on all three retain datasets.
@@
As we will demonstrate in Section 5.1, there is generally a trade-off between forget performance and retain performance. For example, a model can have a near-zero ROUGE score on the forget set but is completely unusable if the model always outputs gibberish given any prompt. Hence, we need to determine a target forget performance as a stopping criterion to facilitate direct comparison between different unlearning methods. In our experiments, we use the ROUGE score of the retraining baseline on the forget set as the stopping criterion, since retraining corresponds to an ideal scenario where the model has never been exposed to the forget set  [^1] By definition the retraining baseline has a forget quality of 1.0, representing an upper bound for this metric. Hence, it is infeasible to use forget quality as the stopping criterion for forget performance.. Following Yao et al. (2024), we match all models to the target ROUGE score by adjusting the learning rate.
@@
In addition to TOFU, we assess if the unlearned model preserves general utilities on well-established benchmarks, including ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021) and OpenBookQA Mihaylov et al. (2018). Since solving these general tasks does not involve knowledge about the data we aim to unlearn, the model after unlearning should maintain as much performance as possible. We follow the default evaluation setting from Gao et al. (2023) and report accuracy on all four tasks under the zero-shot setting.